# LATex: Leveraging Attribute-based Text Knowledge for Aerial-Ground Person Re-Identification

### v2

@@ Abstract
As an important task in intelligent transportation systems, Aerial-Ground person Re-IDentification (AG-ReID) aims to retrieve specific persons across heterogeneous cameras in different viewpoints. Previous methods typically adopt deep learning-based models, focusing on extracting view-invariant features. However, they usually overlook the semantic information in person attributes. In addition, existing training strategies often rely on full fine-tuning large-scale models, which significantly increases training costs. To address these issues, we propose a novel framework named LATex for AG-ReID, which adopts prompt-tuning strategies to leverage attribute-based text knowledge. More specifically, we first introduce the Contrastive Language-Image Pre-training (CLIP) model as the backbone, and propose an Attribute-aware Image Encoder (AIE) to extract both global semantic features and attribute-aware features from input images. Then, with these features, we propose a Prompted Attribute Classifier Group (PACG) to predict person attributes and obtain attribute representations. Finally, we design a Coupled Prompt Template (CPT) to transform attribute representations and view information into structured sentences. These sentences are processed by the text encoder of CLIP to generate more discriminative features. As a result, our framework can fully leverage attribute-based text knowledge to improve AG-ReID performance. Extensive experiments on three AG-ReID benchmarks demonstrate the effectiveness of our proposed methods. The source code will be available.
Index Terms:
Aerial-Ground Person Re-identification, Image-Text Retrieval, Attribute Prediction, Prompt Learning.
@@ I Introduction
Person Re-IDentification (ReID) aims to retrieve the same individual across different cameras. In recent years, ReID has attracted considerable interest [1, 2, 3, 4, 5, 6, 7, 8] due to its wide range of applications, including intelligent surveillance and transportation system. More recently, ReID across heterogeneous camera viewpoints, especially Aerial-Ground person ReID (AG-ReID), has become a more realistic transportation scenario [9, 10, 11] due to the development of drones and advancements in aerial surveillance. Unlike traditional ReID tasks [12, 13, 14], AG-ReID amplifies the challenges posed by viewpoint variations due to drastic changes between different cameras. These variations significantly affect the distribution of body parts [15], making it more difficult to learn robust features that remain consistent across diverse views. Therefore, previous methods [9, 11] focus on mitigating the negative effects of drastic view changes and learning viewpoint-robust image features. However, they often overlook the potential of leveraging person attribute information. As shown in Fig. 1(a) and Fig. 1(c), different camera viewpoints may result in significant visual differences. Despite these significant visual differences, person attributes such as ethnicity, gender, and clothing remain unaffected. This stability provides consistent information to obtain robust cross-view features. Meanwhile, existing methods [9, 11] rely on the full fine-tuning strategy, significantly raising training costs. Fortunately, prompt-tuning [16, 17, 18, 19] offers a way to reduce the training cost. It also effectively integrates the pre-trained knowledge of large-scale models into specific domains [20, 21].
@@
Motivated by the above observations, we propose a novel framework named LATex for AG-ReID, which leverages attribute-based text knowledge via prompt-tuning strategies to enhance the feature discrimination. Specifically, our framework consists of three key components: an Attribute-aware Image Encoder (AIE), a Prompted Attribute Classifier Group (PACG), and a Coupled Prompt Template (CPT). First, we introduce AIE to fine-tune the Contrastive Language-Image Pre-training (CLIP) model [22] with learnable prompts, transferring the powerful pre-trained knowledge to AG-ReID. In addition, AIE incorporates attribute tokens to enable fine-grained perception of person attribute information. Then, PACG is employed to further enhance AIE’s attribute perception capabilities and generate attribute representations. Afterwards, CPT is proposed to transform attribute representations and view information into structured sentences. Finally, these sentences are processed by CLIP’s text encoder, enabling more accurate person retrieval across different camera viewpoints by explicitly leveraging information hidden in the attributes. Extensive experiments on three AG-ReID benchmarks fully validate the effectiveness of our proposed framework.
@@
  • New insight. We observe the distinct benefits of person attribute information for AG-ReID tasks. This insight inspires us to consider the problem from an attribute-based perspective. Based on the attribute consistency, we introduce a practical method to mitigate the challenges in AG-ReID posed by drastic viewpoint changes.
@@ II-A View-Homogeneous ReID
Person ReID is a long-standing task in computer vision and machine learning, gaining significant attention due to its wide range of real-world applications [23, 24, 25, 26, 27]. Previous research has primarily focused on view-homogeneous scenarios, where all cameras in the surveillance network are assumed to operate under the similar viewpoint. Coarsely, the view-homogeneous ReID can be categorized into two types: ground-view and aerial-view. In fact, the ground-view ReID has been widely researched with the support of various datasets, such as Market1501 [28], MSMT17 [29] and CUHK03 [30]. As a consequence, notable advancements have been achieved, primarily categorized into CNN-based methods [31, 32, 33] and Transformer-based methods [34, 35, 36]. For CNN-based methods, Sun et al. [37] and Wang et al. [38] enhance global feature representations by dividing the person image into several parts and extracting the part-level features. Furthermore, Luo et al. [39] provide a strong ReID baseline by introducing some useful tricks. Focusing on the computational efficiency, Quan et al. [40] successfully construct a compact model, namely Auto-ReID, to obtain local discriminative features. In recent years, many methods based on Vision Transformer (ViT) have emerged in the ReID community. For example, He et al. [41] first introduce Transformer into person ReID, achieving promising results. Afterwards, many researchers further leverage Transformers to extract more discriminative person representations [42, 7, 36, 43]. Beside the spatial cues, Li et al. [44] extract high-frequency information of person images to obtain stable representations for ReID. In terms of aerial-view ReID, UAV-Human [45] and PRAI-1581 [46] are the primary benchmarks in this person ReID sub-task. As for advanced methods, Qiu et al. [47] introduce a key-point disentangling strategy for aerial-view ReID. To address the challenge of person rotation, Wang et al. [48] propose a rotation exploration for aerial-view ReID. Although effective, these approaches perform poorly under drastic viewpoint changes, which inevitably appear in AG-ReID.
@@ II-B Aerial-Ground ReID
Recently, advancements in Unmanned Aerial Vehicle (UAV) technologies have made it feasible to deploy dynamic cameras, enhancing surveillance coverage in regions with sparse ground camera networks. However, it poses significant challenges due to the substantial viewpoint variations between UAV cameras and fixed ground cameras. As a result, directly transferring previous view-homogeneous ReID methods often leads to suboptimal performance. To address this issue, AG-ReID has been proposed as a new sub-task of person ReID. To achieve the model training and evaluation, Nguyen et al. [10] first collect an outdoor scene dataset with person attribute annotations, namely AG-ReID.v1. Afterwards, they extend the AG-ReID.v1 with more identities and viewpoints, as AG-ReID.v2 [49]. Zhang et al. [11] construct a large-scale synthesized AG-ReID benchmark, named CARGO. Recently, Zhang et al. [50] consider video-based AG-ReID and contribute the first benchmark, named G2A-VReID. Then, Nguyen et al. [51] further contribute a large-scale video-based AG-ReID benchmark, named AG-VPReID. As for technical methods, Nguyen et al. [10, 49] propose multi-stream frameworks and use person attributes as auxiliary labels for supervision. Based on ViT, Zhang et al. [11] separate identity-related features from viewpoint-specific features by employing view tokens and an orthogonal loss. Moreover, Wang et al. [52] consider the person local features and introduce a prompt-based framework for better AG-ReID. On the other hand, Wang et al. [53] employ a dynamic token selection strategy to focus on key person regions. Hu et al. [54] propose a view-aware stable diffusion to extract comprehensive features for AG-ReID. Although effective, these methods ignore the benefits of explicitly using person attributes for cross-view retrieval. Based on the observations in Fig. 1, we notice that person attributes remain robust in complex scenarios, providing valuable information for discriminative features. Motivated by this, we fully exploit this advantage to alleviate the viewpoint changes in AG-ReID tasks. Technically, we propose a new feature learning framework named LATex that predicts and leverages attribute knowledge to achieve better performance with fewer trainable parameters.
@@ II-C Prompt-Tuning in Person ReID
Prompt-tuning aims to transfer the knowledge of pre-trained models to unseen domains via trainable prompts. Moreover, it typically requires fewer computation resources than full fine-tuning, while also achieving superior performance. This property makes prompt-tuning widely applicable across various tasks [16, 20, 18]. In the ReID field, Li et al. [20] first exploit vision-language models with prompt-tuning to address the lack of missing concrete text labels in image-based person Re-ID. Li et al. [55] propose prototypical contrastive learning to fine-tune CLIP for object ReID. Wu et al. [56] enhance visible-infrared person ReID with modality-aware and instance-aware visual prompt learning. Wang et al. [57] introduce diverse prompt-tuning methods to distill CLIP for learning discriminative person shape representations. Li et al. [58] propose person prompts and clothes prompts to learn cloth-agnostic features for cloth-changing person ReID. Very recently, Wang et al. [52] introduce self-calibrating and adaptive prompts for AG-ReID. Different from previous works, we not only employ learnable prompts as additional tokens to help pre-trained large models generalize to ReID domains, but also leverage the prompt knowledge to further enhance discriminative representations for AG-ReID.
@@ III-B Attribute-aware Image Encoder
To transfer the rich knowledge of CLIP to the ReID task and extract attribute information, we propose the Attribute-aware Image Encoder (AIE). Previous Transformer-based approaches [9, 11] typically rely on full fine-tuning strategies, which lead to very high training costs. To address this issue, we adopt prompt-tuning strategies to extract discriminative features with reduced training resource requirements. Formally, given the input image $\mathcal{V}\in\mathbb{R}^{H\times W\times 3}$ from different views, we embed $\mathcal{V}$ to obtain the visual embedding $F^{0}=[F_{v}^{0},f_{v}^{0}]$, where $F_{v}^{0}\in\mathbb{R}^{C}$ is the class token and $f_{v}^{0}\in\mathbb{R}^{N\times C}$ are patch tokens. Here, $C$ is the dimension of the token embeddings while $N$ is the total number of patches. Then, for the $i$-th Transformer layer $\varOmega_{i}$, we denote $P^{i}_{v}\in\mathbb{R}^{T\times C}$ as the learnable prompts, i.e., $P^{i}_{v}=\{P^{i,1}_{v},\cdots,P^{i,T}_{v}\}$. Here, $T$ is the total number of person attributes. The learnable prompts $P^{i}_{v}$ are concatenated with $F^{i}$, enabling the perception of attribute-specific information, and then passed through $\varOmega_{i}$ as follows:
@@
Here, $\cdot$] means the concatenation operation along the token dimension. Finally, we extract the attribute prompts $\tilde{P}^{L}_{v}\in\mathbb{R}^{T\times C}$ and the visual class token $F_{v}^{L}\in\mathbb{R}^{C}$ from the final layer of AIE for further processing.
@@ III-C Prompted Attribute Classifier Group
where $\tilde{P}^{L,t}\in\mathbb{R}^{C}$ is the $t$-th attribute prompt of $\tilde{P}_{v}^{L}$. To further enhance the attribute-aware features, we utilize the interdependencies of different person attributes. Formally, we denote other attribute prompts as $\hat{P}^{L}_{v}=\{P_{v}^{L,1},\cdots,P_{v}^{L,t-1},P_{v}^{L,t+1},\cdots,P_{v}^{L,T}\}$. Then, we can obtain interacted feature as follows:
@@
where $\Theta(\cdot)$ is the multi-head cross attention [59]. $Q\in\mathbb{R}^{C}$, $K\in\mathbb{R}^{C}$ and $V\in\mathbb{R}^{C}$ are generated by the corresponding projection matrix $W_{q}$, $W_{k}$ and $W_{v}$, respectively. $\delta$ is the Softmax function. The residual structure enables the model to process input features more flexibly [60]. As observed in daily life, certain person attributes exhibit strong correlations (e.g., gender and clothing), while others show weaker correlations (e.g., height and weight) or are nearly independent (e.g., ethnicity and gender). Thus, we employ a residual-based Feed-Forward Network (FFN) [59] to handle features obtained from two perspectives: direct prediction and attribute dependency-based prediction. This design allows the FFN to adapt to diverse scenarios by effectively capturing attribute correlations, if they exist, while avoiding excessive noises and additional computational overhead caused by irrelevant attribute pairs:
@@
where $\Psi$ is a Multi-Layer Perceptron (MLP). The above attribute-based text representations are then input into CLIP’s text encoder to enhance feature discrimination.
@@ III-D Coupled Prompt Template
Recently, large-scale vision-language models [61], such as CLIP have delivered outstanding performance in many computer vision and natural language processing tasks. As a fundamental component, text templates play an important role in text-based person ReID. For example, Li et al. [20] utilize identity-specific prompt tokens to form a text template, i.e., “A photo of a [learnable tokens] person.” However, this kind of templates learn person attributes implicitly, lacking the explicit supervision and ignoring helpful view information.
To address the above issues, we propose a Coupled Prompt Template (CPT), which is presented as “A [view token] view photo of a [shared tokens] [attribute tokens] person.” This template not only couples identity-independent and attribute-aware information, but also leverages comprehensive knowledge of visual-language models. More specifically, the view token $V$ is an instance-level text token, which depends on the view of person images. When the image is captured by a ground camera, the view token is “CCTV”. When the image is from an aerial camera, it is assigned to “UAV”. In addition, we formulate shared tokens as $M_{1},M_{2},\cdots,M_{K}]$”, where $K$ is the total number of tokens. These instance-shared tokens serve as register tokens [62] to enhance semantic feature representations. As for attribute tokens, we denote them as $A_{1},A_{2},\cdots,A_{T}]$”, which are obtained by PACG and have rich attribute information. By effectively utilizing these diverse tokens to form the structured sentence $\mathcal{S}$, our framework can fully leverage the useful information from person attributes. Finally, we feed the sentence $\mathcal{S}$ to the text encoder $\mathcal{T}(\cdot)$ of CLIP to obtain the text feature $F_{d}\in\mathbb{R}^{C}$:

### v3

@@ Abstract
As an important task in intelligent transportation systems, Aerial-Ground person Re-IDentification (AG-ReID) aims to retrieve specific persons across heterogeneous cameras in different viewpoints. Previous methods typically adopt deep learning-based models, focusing on extracting view-invariant features. However, they usually overlook the semantic information in person attributes. In addition, existing training strategies often rely on full fine-tuning large-scale models, which significantly increases training costs. To address these issues, we propose a novel framework named LATex for AG-ReID, which adopts prompt-tuning strategies to leverage attribute-based text knowledge. Specifically, with the Contrastive Language-Image Pre-training (CLIP) model, we first propose an Attribute-aware Image Encoder (AIE) to extract both global semantic features and attribute-aware features from input images. Then, with these features, we propose a Prompted Attribute Classifier Group (PACG) to predict person attributes and obtain attribute representations. Finally, we design a Coupled Prompt Template (CPT) to transform attribute representations and view information into structured sentences. These sentences are processed by the text encoder of CLIP to generate more discriminative features. As a result, our framework can fully leverage attribute-based text knowledge to improve AG-ReID performance. Extensive experiments on three AG-ReID benchmarks demonstrate the effectiveness of our proposed methods. The source code is available at https://github.com/kevinhu314/LATex.
@@ I Introduction
Person Re-IDentification (ReID) aims to retrieve the same individual across different cameras. In recent years, ReID has attracted considerable interest [1, 2, 3, 4, 5, 6] due to its wide range of applications, including intelligent surveillance and transportation system. More recently, ReID across heterogeneous camera viewpoints, especially Aerial-Ground person ReID (AG-ReID), has become a more realistic application [7, 8, 9] due to the development of drones and advancements in aerial surveillance. In practice, AG-ReID greatly helps traffic wardens to manage large transit hubs and address traffic incidents. Unlike traditional ReID tasks [10, 11], AG-ReID amplifies the challenges posed by viewpoint variations due to drastic changes between different cameras. These variations significantly affect the distribution of body parts, making it more difficult to learn visual features that remain consistent across diverse views. Therefore, previous methods [7, 9] focus on mitigating the negative effects of drastic view changes and learning viewpoint-robust image features. However, they often overlook the potential of leveraging person attributes. As shown in Fig. 1(a) and Fig. 1(c), different camera viewpoints may result in significant visual differences. Despite these significant visual differences, person attributes such as ethnicity, gender, and clothing remain unaffected. This stability provides consistent information to obtain robust cross-view features. Meanwhile, existing methods [7, 9] rely on the full fine-tuning strategy, significantly raising training costs. Fortunately, prompt-tuning [12, 13] offers a way to reduce the training cost. It also effectively integrates the pre-trained knowledge of large-scale models into specific domains [14, 15].
@@
Motivated by the above observations, we propose a novel framework named LATex for AG-ReID, which leverages attribute-based text knowledge via prompt-tuning strategies to enhance the feature discrimination. More specifically, our framework consists of three key components: an Attribute-aware Image Encoder (AIE), a Prompted Attribute Classifier Group (PACG), and a Coupled Prompt Template (CPT). First, we introduce AIE to fine-tune the Contrastive Language-Image Pre-training (CLIP) model [16] with learnable prompts, transferring the powerful pre-trained knowledge to AG-ReID. In addition, AIE incorporates attribute tokens to enable fine-grained perception of person attribute information. Then, PACG is employed to further enhance AIE’s attribute perception capabilities and generate attribute representations. Afterwards, CPT is proposed to transform attribute representations and view information into structured sentences. Finally, these sentences are processed by CLIP’s text encoder, enabling more accurate person retrieval across different camera viewpoints by explicitly leveraging information hidden in the attributes. Extensive experiments on three AG-ReID benchmarks fully validate the effectiveness of our proposed framework.
@@
  • New insight. We observe the distinct benefits of person attributes for AG-ReID tasks. This insight inspires us to consider the problem from an attribute-based perspective. Based on the attribute consistency, we introduce a practical method to mitigate the challenges in AG-ReID posed by drastic viewpoint changes.
@@ II-A View-Homogeneous ReID
Person ReID is a long-standing task in computer vision and machine learning, drawing significant attention due to its wide range of real-world applications [17, 18, 19, 20, 21]. Previous research has primarily focused on view-homogeneous scenarios, where all cameras in the surveillance network are assumed to operate under the similar viewpoint. Coarsely, the view-homogeneous ReID can be categorized into two types: ground-view and aerial-view. In fact, the ground-view ReID has been widely researched with the support of various datasets, such as Market1501 [22], MSMT17 [23] and CUHK03 [24]. As a consequence, notable advancements have been achieved, primarily categorized into CNN-based methods and Transformer-based methods. For CNN-based methods, Sun et al. [25] and Wang et al. [26] enhance global feature representations by dividing the person image into several parts and extracting part-level features. Furthermore, Luo et al. [27] provide a strong ReID baseline by introducing some useful tricks. Focusing on the computational efficiency, Quan et al. [28] successfully construct a compact model, namely Auto-ReID, to obtain local discriminative features. In recent years, many methods based on Vision Transformer (ViT) have emerged in the ReID community. For example, He et al. [29] first introduce Transformer into person ReID, achieving promising results. Afterwards, many researchers further leverage Transformers to extract more discriminative person representations [30, 5, 31, 32, 33]. Beside the spatial cues, Li et al. [34] extract high-frequency information of person images to obtain robust representations for ReID. In terms of aerial-view ReID, UAV-Human [35] and PRAI-1581 [36] are the primary benchmarks. As for advanced methods, Qiu et al. [37] introduce a key-point disentangling strategy for aerial-view ReID. To address the challenge of person rotation, Wang et al. [38] propose a rotation exploration for aerial-view ReID. However, these methods perform poorly under drastic viewpoint changes, which inevitably appear in AG-ReID.
@@ II-B Aerial-Ground ReID
Recently, advancements in Unmanned Aerial Vehicle (UAV) technologies have made it feasible to deploy dynamic cameras, enhancing surveillance coverage in regions with sparse ground camera networks. However, it poses significant challenges due to the substantial viewpoint variations between UAV cameras and fixed ground cameras. As a result, directly transferring previous view-homogeneous ReID methods often leads to suboptimal performance. To address this issue, AG-ReID has been proposed as a new sub-task of person ReID. To achieve the model training and evaluation, Nguyen et al. [8] collect an outdoor scene dataset with person attribute annotations, namely AG-ReID.v1. Afterwards, they extend the AG-ReID.v1 with more identities and viewpoints, as AG-ReID.v2 [39]. Zhang et al. [9] construct a large-scale synthesized AG-ReID benchmark, named CARGO. Recently, Zhang et al. [40] consider video-based AG-ReID and contribute the first benchmark, named G2A-VReID. Then, Nguyen et al. [41] further contribute a large-scale video-based AG-ReID benchmark, named AG-VPReID. As for technical methods, Nguyen et al. [8, 39] propose multi-stream frameworks and use person attributes as auxiliary labels for supervision. Based on ViT, Zhang et al. [9] separate identity-related features from viewpoint-specific features by employing view tokens and an orthogonal loss. Moreover, Wang et al. [42] consider the person local features and introduce a prompt-based framework for better AG-ReID. On the other hand, Wang et al. [43] employ a dynamic token selection strategy to focus on key person regions. Although effective, these methods ignore the benefits of explicitly using person attributes for cross-view retrieval. Based on the observations in Fig. 1, we notice that person attributes remain robust in complex scenarios, providing valuable information for discriminative features. Motivated by this, we fully exploit this advantage to alleviate the viewpoint changes in AG-ReID tasks. Technically, we propose a new feature learning framework named LATex that predicts and leverages attribute knowledge to achieve better performance with fewer trainable parameters.
@@ II-C Prompt-Tuning in Person ReID
Prompt-tuning aims to transfer the knowledge of pre-trained models to unseen domains via trainable prompts. Moreover, it typically requires fewer computation resources than full fine-tuning, while also achieving superior performance. This property makes prompt-tuning widely applicable across various tasks [12, 14, 13]. In the ReID field, Li et al. [14] first exploit vision-language models with prompt-tuning to address the lack of missing concrete text labels in image-based person Re-ID. Yu et al. [44] further propose text-free CLIP with prompt-tuning for video-based person ReID. Wu et al. [45] enhance visible-infrared person ReID with modality-aware and instance-aware visual prompt learning. Wang et al. [46] introduce diverse prompt-tuning methods to distill CLIP for learning discriminative person shape representations. Li et al. [47] propose person prompts and clothes prompts to learn cloth-agnostic features for cloth-changing person ReID. Very recently, Wang et al. [42] introduce self-calibrating and adaptive prompts for AG-ReID. Yu et al. [48] propose a hybrid CLIP-Mamba framework for person ReID. Wang et al. [49] adopt attribute prompt composition for object ReID. Different from previous works, we not only employ learnable prompts as additional tokens to help pre-trained large models generalize to ReID domains, but also leverage prompt knowledge to further enhance discriminative features for AG-ReID.
@@ III-B Attribute-aware Image Encoder
To transfer the rich knowledge of CLIP to the ReID task and extract attribute information, we propose the Attribute-aware Image Encoder (AIE). Previous Transformer-based approaches [7, 9] typically rely on full fine-tuning strategies, which lead to very high training costs. To address this issue, we adopt prompt-tuning strategies to extract discriminative features with reduced training resource requirements. Formally, given the input image $\mathcal{V}\in\mathbb{R}^{H\times W\times 3}$ from different views, we embed $\mathcal{V}$ to obtain the visual embedding $F^{0}=[F_{v}^{0},f_{v}^{0}]$, where $F_{v}^{0}\in\mathbb{R}^{C}$ is the class token and $f_{v}^{0}\in\mathbb{R}^{N\times C}$ are patch tokens. Here, $C$ is the dimension of the token embeddings while $N$ is the total number of patches. Then, for the $i$-th Transformer layer $\varOmega_{i}$, we denote $P^{i}_{v}\in\mathbb{R}^{\hat{T}\times C}$ as the learnable prompts, i.e., $P^{i}_{v}=\{P^{i,1}_{v},\cdots,P^{i,\hat{T}}_{v}\}$. Here, the first $T$ prompts are treated as attribute-aware prompts. The remaining $\hat{T}-T$ prompts are employed to support the fine-tuning of AIE. The learnable prompts $P^{i}_{v}$ are concatenated with $F^{i}$, enabling the perception of attribute-specific information, and then passed through $\varOmega_{i}$ as follows:
@@
Here, $\cdot$] means the concatenation operation along the token dimension. Finally, we extract attribute prompts $\tilde{P}^{L}_{v}\in\mathbb{R}^{T\times C}$ and the class token $F_{v}^{L}\in\mathbb{R}^{C}$ from the final layer of AIE for further processing.
@@ III-C Prompted Attribute Classifier Group
where $\tilde{P}^{L,t}\in\mathbb{R}^{C}$ is the $t$-th attribute prompt of $\tilde{P}_{v}^{L}$. To further enhance attribute-aware features, we utilize the interdependencies of different person attributes. Formally, we denote other attribute prompts as $\hat{P}^{L}_{v}=\{P_{v}^{L,1},\cdots,P_{v}^{L,t-1},P_{v}^{L,t+1},\cdots,P_{v}^{L,T}\}$. Then, we can obtain interacted feature as follows:
@@
where $\Theta(\cdot)$ is the multi-head cross attention [50]. $Q\in\mathbb{R}^{C}$, $K\in\mathbb{R}^{C}$ and $V\in\mathbb{R}^{C}$ are generated by the corresponding projection matrix $W_{q}$, $W_{k}$ and $W_{v}$, respectively. $\delta$ is the Softmax function. The residual structure enables the model to process input features more flexibly [51]. As observed in daily life, certain person attributes exhibit strong correlations (e.g., gender and clothing), while others show weaker correlations (e.g., height and weight) or are nearly independent (e.g., ethnicity and gender). Thus, we employ a residual-based Feed-Forward Network (FFN) to handle features obtained from two perspectives: direct prediction and attribute dependency-based prediction. This design allows the FFN to adapt to diverse scenarios by effectively capturing attribute correlations, if they exist, while avoiding excessive noises and additional computational overhead caused by irrelevant attribute pairs:
@@
where $\Psi$ is a Multi-Layer Perceptron (MLP). The resulting representations serve as continuous textual tokens, which are fed into CLIP’s text encoder to enhance feature discrimination.
@@ III-D Coupled Prompt Template
Recently, large-scale vision-language models have delivered outstanding performance in many computer vision and natural language processing tasks. As a fundamental component, text templates play an important role in text-based person ReID. For example, Li et al. [14] utilize identity-specific prompt tokens to form a text template, i.e., “A photo of a [learnable tokens] person.” However, this kind of templates learn person attributes implicitly, lacking the explicit supervision and ignoring helpful view information.
To address the above issues, we propose a Coupled Prompt Template (CPT), which is presented as “A [view token] view photo of a [shared tokens] [attribute tokens] person.” This template not only couples identity-independent and attribute-aware information, but also leverages comprehensive knowledge of visual-language models. More specifically, the view token $V$ is an instance-level text token, which depends on the view of person images. For instance, the view token is designated as “CCTV” for images captured by ground views and “UAV” for those from an aerial view. Functionally, the proposed framework is readily extensible to other viewpoints (e.g., a wearable device view) by simply introducing a new view token, requiring no architectural changes. In addition, we formulate shared tokens as $M_{1},M_{2},\cdots,M_{K}]$”, where $K$ is the total number of tokens. These instance-shared tokens serve as register tokens [52] to enhance semantic feature representations. As for attribute tokens, we denote them as $A_{1},A_{2},\cdots,A_{T}]$”, which are obtained by PACG and have rich attribute information. By effectively utilizing these diverse tokens to form the structured sentence $\mathcal{S}$, our framework can fully leverage the useful information from person attributes. Finally, we feed the sentence $\mathcal{S}$ to the text encoder $\mathcal{T}(\cdot)$ of CLIP to obtain the text feature $F_{d}\in\mathbb{R}^{C}$: